Light travels with speed of 2x10^8  m/s in crown glass of refractive index 1.5  .  What is the speed of light in dense flint  glass of refractive index 1.8 ?

Grade:12

3 Answers

VIRAJ
10 years ago

is the answer 1.67x108.........

mukunda
10 years ago

hi    ......             take it this is audio type solution for your problem and answer is 1.97 x 108 m/s  .  bye..

suhel
10 years ago

light travels with speed va = 2x10^8m/s

glass refractive index na = 1.5

flint glass refractive index nb = 1.8

To find :-

        speed of light ij dense flint glass vb

                      na/nb = vb/va

                     1.5/1.8 =vb/2x10^8m/s

                     0.8333x2x10^8 = vb

                     vb=1.6666x10^8

        therefore, the speed of light in dense flint glass is 1.666x10^8m/s